# On the Geometry of Deep Learning

### v1

@@ Introduction
Despite this exciting empirical progress, however, the precise mechanisms by which deep learning works so well remain relatively poorly understood, adding an air of mystery to the entire field. Ongoing attempts to build a rigorous mathematical framework have been stymied by the fact that, while deep networks are locally simple, they are globally complicated. Hence, they have primarily been studied as “black boxes” and mainly empirically. This approach greatly complicates analysis to understand the both the success and failure modes of deep networks. This approach also greatly complicates deep learning system design, which today proceeds alchemistically rather than from rigorous design principles. And this approach greatly complicates addressing higher level issues like trustworthiness (can we trust a black box?), sustainability (ever-growing computations lead to a growing environmental footprint), and social responsibility (fairness, bias, and beyond).
In this paper, we overview one promising avenue of progress at the mathematical foundation of deep learning: the connection between deep networks and function approximation by affine splines (continuous piecewise linear functions in multiple dimensions). In particular, we will overview work over the past decade on understanding certain geometrical properties of a deep network’s affine spline mapping, in particular how it tessellates its input space. As we will see, the affine spline connection and geometrical viewpoint provide a powerful portal through which to view, analyze, and improve the inner workings of a deep network.
@@
The state-of-the-art in deep learning is a rapidly moving target, and so we focus on the bedrock of modern deep networks, so-called feedforward neural networks employing piecewise linear activation functions. While our analysis does not fully cover some very recent methods, most notably transformer networks, the networks we study are employed therein as key building blocks. Moreover, since we focus on the affine spline viewpoint, we will not have the opportunity to discuss other interesting geometric work in deep learning, including tropical geometry [zhang2018tropical] and beyond. Finally, to spin a consistent story line, we will focus primarily on work from our group; we will however review several key results developed by others. Our bibliography is concise, and so the interested reader is invited to explore the extensive works cited in the papers we reference.
@@
Machine learning in 200 words or less. In supervised machine learning, we are given a collection of $n$ training data pairs $\left\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\right\}_{i=1}^{n}$; $\boldsymbol{x}_{i}$ is termed the data and $\boldsymbol{y}_{i}$ the label. Without loss of generality, we will take $\boldsymbol{x}_{i}\in{\mathbb{R}}^{D},\boldsymbol{y}_{i}\in{\mathbb{R}}^{C}$ to be column vectors, but in practice they are often tensors.
@@
One solves the prediction problem by first designing a parameterized model $f_{\Theta}$ with parameters $\Theta$ and then learning or training by optimizing $\Theta$ to make $f_{\Theta}(\boldsymbol{x}_{i})$ as close as possible to $\boldsymbol{y}_{i}$ on average in terms of some distance or loss function ${\cal L}$, which is often called the training error.
@@
Here $\Theta$ is the collection of parameters from each layer, $\theta^{(\ell)}$, $\ell=1,\dots,L$. We will omit the parameters $\Theta$ or $\theta^{(\ell)}$ from our notation except where they are critical, since they are ever-present in the discussion below.
The $\ell$-th deep network layer $f^{(\ell)}$ takes as input the vector $\boldsymbol{z}^{(\ell-1)}$ and outputs the vector $\boldsymbol{z}^{(\ell)}$ by combining two simple operations
@@
The parameters $\theta^{(\ell)}$ of the layer are the elements of the weight matrix $\boldsymbol{W}^{(\ell)}$ and the bias vector $\boldsymbol{b}^{(\ell)}$. Special network structures have been developed to reduce the generally quadratic cost of multiplying by the $\boldsymbol{W}^{(\ell)}$. One notable class of networks constraints $\boldsymbol{W}^{(\ell)}$ to be a circulant matrix, so that $\boldsymbol{W}^{(\ell)}\boldsymbol{z}^{(\ell)}$ corresponds to a convolution, giving rise to the term ConvNet for such models. Even with this simplification, it is common these days to work with networks with billions of parameters.
@@
Throughout this paper, we focus on networks that use this activation, although the results hold for any continuous piecewise linear nonlinearity (e.g., absolute value $\sigma(u)=|u|$). Special activations are often employed at the last layer $f^{(L)}$, from the linear activation $\sigma(u)=u$ to the softmax that converts a vector to a probability histogram. These activations do not affect our analysis below. It is worth point out, but beyond the scope of this paper, that it is possible to generalize the results we review below to a much larger class of smooth activation functions (e.g., sigmoid gated linear units, Swish activation) by adopting a probabilistic viewpoint [balestriero2018hard].
@@
Figure 1: A 6-layer deep network. Purple, blue, and yellow nodes represent the input, neurons, and output, respectively. The width of layer 2 is 5, for example. The links between the nodes represent the elements of the weight matrices $\boldsymbol{W}^{(\ell)}$. The sum with the bias $\boldsymbol{b}^{(\ell)}$ and subsequent activation $\sigma(\cdot)$ are implicitly performed at each neuron.
Learning. To learn to fit the training data with a deep network, we tune the parameters $\boldsymbol{W}^{(\ell)},\boldsymbol{b}^{(\ell)},\ell=1,\dots,L$ such that, on average, when training datum $\boldsymbol{x}_{i}$ is input to the network, the output $\widehat{\boldsymbol{y}_{i}}=f(\boldsymbol{x}_{i})$ is close to $\boldsymbol{y}_{i}$ as measured by some loss function ${\cal L}$. Two loss functions are ubiquitous in deep learning; the first is the classical squared error based on the two-norm
@@
The other is the cross-entropy which is oft-used in classification tasks.
@@
Standard learning practice is to use some flavor of gradient (steepest) descent that iteratively reduces ${\cal L}$ by updating the parameters $\boldsymbol{W}^{(\ell)},\boldsymbol{b}^{(\ell)}$ by subtracting a small scalar multiple of the partial derivatives of ${\cal L}$ with respect to those parameters.
@@
In practice, since the number of training data pairs $n$ can be enormous, one calculates the gradient of ${\cal L}$ for each iteration using only a subset of data points and labels called a minibatch.
@@
Note that even a nice loss function like (4) will have a multitude of local minima due to the nonlinear activation at each layer coupled with the composition of multiple such layers. Consequently, numerous heuristics have been developed to help navigate to high-performing local minima. In modern deep networks, the number of neurons is usually so enormous that, by suitably optimizing the parameters, one can nearly interpolate the training data. (We often drop the “nearly” below for brevity.) What distinguishes the performance of one deep network architecture from another, then, is what it does away from the training points, i.e., how well it generalizes to unseen data.
@@
Deep nets break out. Despite neural networks existing in some form for over 80 years, their success was limited in practice until the AI boom of 2012. This sudden growth was enabled by three converging factors: i) going deep with many layers, ii) training on enormous data sets, and iii) new computing architectures based on graphics processing units (GPUs).
@@
The spark that ignited the AI boom was the Imagenet Challenge 2012 where teams competed to best classify a set of input images $\boldsymbol{x}$ into one of 1000 categories. The Imagenet training data was about $n=1.3$ million, 150,000-pixel color digital images human-labeled into 1000 classes, such as ‘bird,’ ‘bus,’ ‘sofa.’ 2012 was the first time a deep network won the Challenge; AlexNet, a ConvNet with 62 million parameters in five convolutional layers followed by three fully connected layers achieved and accuracy of 60%. Subsequent competitions featured only deep networks, and, by the final competition in 2017, they had reached 81% accuracy, which is arguably better than most humans can achieve.
@@
Black boxes. Deep networks with dozens of layers and millions of parameters are powerful for fitting and mimicking training data, but also inscrutable. Deep networks are created from such simple transformations (e.g., affine transform and thresholding) that it is maddening the composition of several of them so complicates analysis and defies deep understanding. Consequently, deep learning practitioners tend to treat them as black boxes and proceed empirically using an alchemical development process that focuses primarily on the inputs $\boldsymbol{x}$ and outputs $f(\boldsymbol{x})$ of the network. To truly understand deep networks we need to be able to see inside the black box as a deep network is learning and predicting. In the sequel, we will discuss one promising line of work in this vein that leverages the fact that deep networks are affine spline mappings.
@@
As we will now explain, deep networks are tractable multidimensional extensions of the familiar one-dimensional continuous piecewise linear functions depicted on the left in Figure 2. When such continuous piecewise functions are fit to training data, we refer to them as affine splines for short.
@@
Deep networks implement one particular extension of the affine spline concept to a multidimensional domain and range. As we will see in the next section, a deep network generalizes the intervals of the independent variable over which a piecewise affine function is simply affine (recall Figure 2) to an irregular tessellation (tiling) of the network’s $D$-dimensional input space into convex polytopes. Let $\Omega$ denote the tessellation and $\omega\in\Omega$ an individual tile. (The jargon for the polytope tiles is “linear region” [montufar2014number].)
@@
The weights and biases of the first layer determine not only the tessellation of the input space but also an affine transformation on each tile to implement (5). Explicit formulas for $\boldsymbol{A}_{\omega},\boldsymbol{c}_{\omega}$ are available in [romain]. It should be clear that, since all of the transformations in (6) are continuous, so must be the affine spline (5) corresponding to the first layer.
The tessellation corresponding to the composition of two or more layers follows an interesting subdivision process akin to a “tessellation of tessellations” [romain]. For example, the second layer creates a hyperplane arrangement in its input space, which happens to be the output space of layer one. Thus, these hyperplanes can be pulled back through layer one to its input space by performing the same process as above but on a tile-by-tile basis relative to the layer-one tessellation and its associated affine transforms. The effect on the layer-two hyperplanes is that they are folded each time they cross a hyperplane created by layer 1. Careful inspection of the tessellation in Figure 3 reveals many examples of such hyperplane folding. Similarly, the hyperplanes created by layer three will be folded every time they encounter a hyperplane in the input space from layers one or two.
Much can be said about this folding process, including a formula for the dihedral angle of a folded hyperplane as a function of the network’s weights and biases. However, the formulae for the angles and affine transformations unfortunately become unwieldy for more than two layers. Finding simplifications for these attributes is an interesting open problem as are the connections to other subdivision processes like wavelets and fractals.
The theory of hyperplane arrangements is rich and tells us that, generally speaking, the number of tiles grows rapidly with the number of neurons in each layer. Hence, we can expect even modestly sized deep networks to have an enormous number of tiles in their input space, each with a corresponding affine transformation from input to output space. Importantly, though, the affine transformations are highly coupled because the overall mapping (5) must remain continuous. This means that the class of functions that can be represented using a deep network is considerably smaller than if the mapping could be uncoupled and/or discontinuous. Understanding what deep learning practitioners call the network’s “implicit bias” remains an important open problem.
@@
SplineCam is an exact method for computing and visualizing a deep network’s spline tessellation over a specified low-dimensional region of the input space, typically a bounded two-dimensional planar slice [splinecam]. SplineCam uses an efficient graph data structure to encode the intersections of the hyperplanes (from the various layers) that pass through the slice and then uses a fast heuristic breadth-first search algorithm to identify tiles from the graph. All of the computations besides the search can be vectorized and computed on GPUs to enable the visualization of even industrial-scale deep networks.
Figure 5 depicts a SplineCam slice along the plane defined by three training images for a 5-layer ConvNet trained to classify between Egyptian and Tabby cat photos. The first thing we notice is the extraordinarily large number of tiles in just this small region of the 4096-dimensional input space. It can be shown that the decision boundary separating Egyptian and Tabby cats corresponds to a single hyperplane in the final layer that is folded extensively from being pulled back through the previous four layers [romain]. Photos falling in the lower left of the slice are classified as Tabbies, while photos falling in the lower right are classified as Egyptians. The density of tiles is also not uniform and varies across the input space.
@@
It has been known since the late 1980s that even a two-layer neural network is a universal approximator, meaning that, as the number of neurons grows, one can approximate an arbitrary continuous function over a Borel measurable set to arbitrary precision [cybenko]. But, unfortunately, while two-layer networks are easily capable of interpolating a set of training data, in practice they do a poor job generalizing to data outside of the training set. In contrast, deep networks with $L\gg 2$ layers have proved over the past 15 years that they are capable of both interpolating and generalizing well.
@@
Further work has worked to link the self-similar nature of the tessellation to good approximation. Using self-similarity, one can construct new function spaces for which deeper networks provide better approximation rates (see [devore2021neural, daubechies2022nonlinear] and the references therein). The benefits of depth stem from the fact that the model is able to replicate a part of the function it is trying to approximate in many different places in the input space and with different scalings or orientations. Extending these results, which currently hold only for one-dimensional input and output spaces, to multidimensional signals is an interesting open research avenue.
@@
In particular, we can make a fair, quantitative comparison between the loss landscapes of the ConvNet and ResNet architectures by comparing their singular values. The key finding is that the condition number of a ResNet (the ratio of the largest to smallest singular value) is bounded, while that of the ConvNet is not [rolf]. This means that the local loss landscape of a ResNet with skip connections is provably better conditioned than that of a ConvNet and thus less erratic, less eccentric, and with local minima that are more accommodating to gradient-based optimization.
@@
Astute readers might see a connection to the standard data preprocessing step of data normalization and centering; the main difference is that this processing is performed before each and every gradient learning step. Batch normalization often greatly aids the optimization of a wide variety of deep networks, helping it to find a better (lower) minimum quicker. But the reasons for its efficacy are poorly understood.
We can make progress on understanding batch normalization by again leaning on the affine spline viewpoint. Let’s focus on the effect of batch normalization at initialization just before gradient learning begins; the effect is pronounced and it is then easy to extrapolate regarding what happens at subsequent gradient steps. Prior to learning, a deep network’s weights are initialized with random numbers. This means that the initial hyperplane arrangement is also random.
The key finding of [BN-arxiv] is that batch normalization adapts the geometry of a deep network’s spline tessellation to focus the network’s attention on the training data $\boldsymbol{x}_{i}$. It does this by adjusting the angles and offsets of the hyperplanes that form the boundaries of the polytopal tiles to increase their density in regions of the input space inhabited by the training data, thereby enabling finer approximation there.
@@
Figure 8 provides clear evidence of batch normalization’s adaptive prowess. We initialize an 11-layer deep network with a two-dimensional input space three different ways to train on data with a star-shaped distribution. We plot the density of the hyperplanes (basically, the number of hyperplanes passing through local regions of the input space) created by layers 3, 7, and 11 for three different layer configurations: i) the standard layer (6) with bias $\boldsymbol{b}^{(\ell)}={\bf 0}$; ii) the standard layer (6) with random bias $\boldsymbol{b}^{(\ell)}$; iii) the batch normalization layer (8). In all three cases, the weights $\boldsymbol{W}^{(\ell)}$ were initialized randomly. We can make several observations. First, constraining the bias to be zero forces the network into a central hyperplane arrangement tessellation that is not amenable to aligning with the data. Second, randomizing both the weights and biases splays the tiles over the entire input space, including many places where the training data is not. Third, batch normalization focuses the hyperplanes from all three of the layers onto the regions where the training data lives.
@@
Current deep learning practice treats a deep network as a black box and optimizes its internal parameters (weights and biases) to minimize some end-to-end training error like the squared loss in (4). While this approach has proved mightily successful empirically, it provides no insight into how learning is going on inside the network nor how to improve it. Clearly, as we adjust the parameters to decrease the loss function using gradient descent, the tessellation will change dynamically. Can we use this insight to learn something new about what goes on inside a deep network during learning?
Consider a deep network learning to classify photos of handwritten digits 0–9. Figure 9 deploys SplineCam to visualize a portion of a 2D slice of the input space of the network defined by three data points in the MNIST handwritten digit training dataset [alwaysgrok]. At left, we see that the tessellation at initialization (before we start learning) is in disarray due to the random weights and biases and nonuse of batch normalization (more on this later). The tessellation is random, and the training error is large.
In the middle, we see the tessellation after convergence to near-zero training error, when most of the digits are on the correct side of their respective decision boundaries. Not shown by the figure is the fact that the network also generalizes well to unseen test data at this juncture. High density suggests that even a continuous piecewise affine function can be quite rugged around these points [sebastianParis]. Indeed, the false coloring indicates that the 2-norms of the $\boldsymbol{A}_{\omega}$ matrices has increased around the training images, meaning that their “slopes” have increased. As a consequence, the overall spline mapping $f(\boldsymbol{x})$ is now likely more rugged and more sensitive to changes in the input $\boldsymbol{x}$ as measured by a local (per-tile) Lipschitz constant. In summary, at (near) interpolation, the gradient learning iterations have in some sense accomplished their task (near zero training error) but with elevated sensitivity of $f(\boldsymbol{x})$ to changes in $\boldsymbol{x}$ around the training data points as compared to the random initialization.
@@
Figure 9: SplineCam visualization of a slice of the input space defined by three MNIST digits being classified by a 4-layer MLP of width 200. The false color map (vivirdis) encodes the 2-norm of the $\boldsymbol{A}_{\omega}$ matrix defined on each tile according to purple (low), green (medium), yellow (high). The decision boundary is depicted in red. (Adapted from [alwaysgrok].)

### v2

@@ Introduction
Despite this exciting empirical progress, however, the precise mechanisms by which deep learning works so well remain relatively poorly understood, adding an air of mystery to the entire field. Ongoing attempts to build a rigorous mathematical framework have been stymied by the fact that, while deep networks are locally simple, they are globally complicated. Hence, they have been studied primarily as “black boxes” and mainly empirically. This approach greatly complicates analysis to understand both the success and failure modes of deep networks. This approach also greatly complicates deep learning system design, which today proceeds alchemistically rather than from rigorous design principles. And this approach greatly complicates addressing higher level issues like trustworthiness (can we trust a black box?), sustainability (ever-growing computations lead to a growing environmental footprint), and social responsibility (fairness, bias, and beyond).
In this paper, we overview one promising avenue of progress at the mathematical foundation of deep learning: the connection between deep networks and function approximation by affine splines (continuous piecewise linear functions in multiple dimensions). In particular, we overview work over the past decade on understanding certain geometrical properties of a deep network’s affine spline mapping, in particular how it tessellates its input space. As we will see, the affine spline connection and geometrical viewpoint provide a powerful portal through which to view, analyze, and improve the inner workings of deep networks.
@@
The state-of-the-art in deep learning is a rapidly moving target, and so we focus on the bedrock of modern deep networks, so-called feedforward neural networks employing piecewise linear activation functions. While our analysis does not fully cover some very recent methods, most notably transformer networks, the networks we study are employed therein as key building blocks. Moreover, since we focus on the affine spline viewpoint, we will not have the opportunity to discuss other interesting geometric work in deep learning, including tropical geometry [zhang2018tropical] and beyond. Finally, to spin a consistent story line, we will focus primarily on work from our group; we will, however, review several key results developed by others. Our bibliography is concise, and so we invite the interested reader to explore the extensive works cited in the papers we reference.
@@
Machine learning in 200 words or less. In supervised machine learning, we are given a collection of $n$ training data pairs $\left\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\right\}_{i=1}^{n}$; $\boldsymbol{x}_{i}$ is termed the data and $\boldsymbol{y}_{i}$ the label. Without loss of generality, we take $\boldsymbol{x}_{i}\in{\mathbb{R}}^{D},\boldsymbol{y}_{i}\in{\mathbb{R}}^{C}$ to be column vectors, but in practice they are often tensors.
@@
One solves the prediction problem by first designing a parameterized model $f_{\Theta}$ with parameters $\Theta$ and then learning or training by optimizing $\Theta$ to make $f_{\Theta}(\boldsymbol{x}_{i})$ as close as possible to $\boldsymbol{y}_{i}$ on average in terms of some distance or loss function ${\cal L}(\Theta)$ that measures the training error.
@@
Here $\Theta$ is the collection of parameters from each layer, $\theta^{(\ell)}$, $\ell=1,\dots,L$. We omit the parameters $\Theta$ or $\theta^{(\ell)}$ from our notation except when they are critical, since they are ever-present in the discussion below.
The $\ell$-th deep network layer $f^{(\ell)}$ takes as input the vector $\boldsymbol{z}^{(\ell-1)}$ and outputs the vector $\boldsymbol{z}^{(\ell)}$ by combining two simple operations:
@@
The parameters $\theta^{(\ell)}$ of the layer are the elements of the weight matrix $\boldsymbol{W}^{(\ell)}$ and the bias vector $\boldsymbol{b}^{(\ell)}$. Special network structures have been developed to reduce the generally quadratic cost of multiplying by $\boldsymbol{W}^{(\ell)}$. One notable class of networks constrains $\boldsymbol{W}^{(\ell)}$ to be a circulant matrix, so that $\boldsymbol{W}^{(\ell)}\boldsymbol{z}^{(\ell)}$ corresponds to a convolution, giving rise to the term ConvNet for such models. Even with this simplification, it is common these days to work with networks with many billions of parameters.
@@
Throughout this paper, we focus on networks using this activation, although the results hold for any continuous piecewise linear nonlinearity (e.g., absolute value, $\sigma(u)=|u|$). Special activations are often employed at the last layer $f^{(L)}$, from the linear activation $\sigma(u)=u$ to the softmax that converts a vector to a probability histogram [goodfellow2016deep]. These activations do not affect our analysis below. It is worth pointing out, but beyond the scope of this paper, that the results we review below extend to a much larger class of smooth activation functions (e.g., sigmoid gated linear units, Swish activation) by adopting a probabilistic viewpoint [balestriero2018hard].
@@
Figure 1: A 6-layer deep network. The purple, blue, and yellow nodes represent the input, neurons, and output, respectively, while the edges represent the affine transformation and activation effected by each layer. The width of layer 2 is 5, for example. The links between the nodes represent the elements of the weight matrices $\boldsymbol{W}^{(\ell)}$. The sum with the bias $\boldsymbol{b}^{(\ell)}$ and subsequent activation $\sigma(\cdot)$ are implicitly performed at each neuron.
Learning. To learn to fit the training data with a deep network, we tune the parameters $\boldsymbol{W}^{(\ell)},\boldsymbol{b}^{(\ell)},\ell=1,\dots,L$ such that, on average, when training datum $\boldsymbol{x}_{i}$ is input to the network, the output $\widehat{\boldsymbol{y}_{i}}=f(\boldsymbol{x}_{i})$ is close to $\boldsymbol{y}_{i}$ as measured by the loss function ${\cal L}$. Two loss functions are ubiquitous in deep learning. The first is the classical squared error based on the two-norm
@@
that is used in regression tasks, where the labels $\boldsymbol{y}_{i}$ are real-valued. The other is the cross-entropy that is oft-used in classification tasks, where the labels are discrete.
@@
Standard learning practice is to use some flavor of gradient (steepest) descent to iteratively reduce ${\cal L}$ by updating the parameters $\boldsymbol{W}^{(\ell)},\boldsymbol{b}^{(\ell)}$ by subtracting a small scalar multiple of the partial derivatives of ${\cal L}$ with respect to those parameters.
@@
In practice, since the number of training data pairs $n$ can be enormous, one calculates the gradient of ${\cal L}$ for each iteration using only a subset of training data points and labels called a minibatch.
@@
Note that even a nice loss function like (4) has a multitude of local minima due to the nonlinear activation $\sigma$ in each layer coupled with the composition of multiple such layers. Consequently, numerous heuristics have been developed to help navigate to high-performing local minima.
In modern deep networks, the number of neurons is usually so gigantic that, by suitably optimizing the parameters, one can nearly interpolate the training data. (We often drop the “nearly” below for brevity.) What distinguishes the performance of one deep network architecture from another, then, is what it does away from the training points, i.e., how well it generalizes to unseen data.
@@
Deep nets break out. Despite neural networks existing in some form for over 80 years, their success was limited in practice until the AI boom of 2012. Sudden rapid progress was enabled by three converging factors: i) going deep with many layers (i.e., big $L$), ii) training on enormous data sets (i.e., big $n$), and iii) new computing architectures based on graphics processing units (GPUs).
@@
The spark that ignited the AI boom was the Imagenet Challenge 2012, where teams competed to best classify a set of input digital photographs into one of 1000 categories. The Imagenet training data comprised about $n=1.3$ million, $D=$ 150,000-pixel color digital images human-labeled into $C=1000$ classes, such as ‘bird,’ ‘bus,’ ‘sofa.’ 2012 was the first time a deep network won the Challenge; AlexNet, a ConvNet with 62 million parameters in five convolutional layers followed by three general layers achieved an accuracy of 60%. Subsequent competitions featured only deep networks, and, by the final competition in 2017, they had reached 81% accuracy, which is arguably better than most humans can achieve.
@@
Black boxes. Deep networks with dozens of layers and millions or even billions of parameters are powerful for fitting and mimicking training data but also inscrutable. It is maddening that the mere composition of simple transformations (i.e., affine transforms and thresholding) so complicates analysis and defies detailed understanding. Consequently, deep learning practitioners tend to treat them as black boxes and proceed empirically using an alchemical development process that focuses primarily on the inputs $\boldsymbol{x}$ and outputs $f(\boldsymbol{x})$ of the network. To truly understand deep networks we need to be able to see inside the black box as a deep network is learning and predicting. In the sequel, we discuss one promising line of work in this vein that leverages the fact that deep networks are affine spline mappings.
@@
As we now explain, deep networks are tractable multidimensional extensions of the familiar one-dimensional continuous piecewise linear functions depicted on the left in Figure 2. When such continuous piecewise functions are fit to training data, we refer to them as affine splines for short.
@@
Deep networks implement one particular extension of the affine spline concept to a multidimensional domain and range. As we will see in the next section, a deep network generalizes the intervals of the independent variable over which a piecewise affine function is purely affine (recall Figure 2) to an irregular tessellation (tiling) of the network’s $D$-dimensional input space into convex polytopes. Let $\Omega$ denote the tessellation and $\omega\in\Omega$ an individual tile. (The deep learning jargon for the polytope tiles is “linear regions” [montufar2014number].)
@@
One may wonder why we would set up this indirect deep network machinery just to implement an affine spline. The reason is that direct spline representations are entirely impractical in machine learning settings, for two reasons. First, we want the polytope tile boundaries to be unconstrained, and even for $D=1$, such “free-knot” splines are combinatorially complex to optimize. Second, it is not clear how to extend the idea of a free-knot spline to even $D=2$, let alone to high dimensions.
@@
The weights and biases of the first layer determine not only the tessellation of the input space but also an affine transformation on each tile to implement (5). Explicit formulas for $\boldsymbol{A}_{\omega},\boldsymbol{c}_{\omega}$ are available in [romain]. It should be clear that, since all of the transformations in (6) are continuous, so must be the affine spline (5) corresponding to the layer.
The tessellation corresponding to the composition of two or more layers follows an interesting subdivision process akin to a “tessellation of tessellations” [romain]. For example, the second layer creates a hyperplane arrangement in its input space, which happens to be the output space of layer one. These hyperplanes can be pulled back through layer one to its input space by performing the same process as above but on a tile-by-tile basis relative to the layer-one tessellation and its associated affine transforms. The effect on the layer-two hyperplanes is that they are folded each time they cross a hyperplane created by layer 1. Careful inspection of the tessellation in Figure 3 reveals many examples of such hyperplane folding. Similarly, the hyperplanes created by layer three are folded every time they encounter a hyperplane in the input space from layers one or two.
Much can be said about this folding process, including a formula for the dihedral angle of a folded hyperplane as a function of the network’s weights and biases [romain]. Interestingly, the collection of dihedral angles of the folded hyperplanes from the final layer of a classification network determines the smoothness of the network’s decision boundaries that partition the input space into regions corresponding to the $C$ classes. Unfortunately, the formulae for the angles and affine transformations become unwieldy for more than two layers. Finding simplifications for these attributes is an interesting open problem as are the connections to other subdivision processes like wavelets and fractals (more on this below).
The theory of hyperplane arrangements is rich and tells us that, generally speaking, the number of tiles grows rapidly with the number of neurons in each layer. Hence, we can expect even modestly sized deep networks to have an enormous number of tiles in their input space, each with a corresponding affine transformation from the input to output space. Importantly, though, the affine transformations are highly coupled because the overall mapping (5) must remain continuous. This means that the class of functions that can be represented using a deep network is considerably smaller than if the mapping could be uncoupled and/or discontinuous. Understanding what deep learning practitioners call the network’s “implicit bias” remains an important open problem.
@@
SplineCam is an exact method for computing and visualizing a deep network’s spline tessellation over a specified low-dimensional region of the input space, typically a bounded two-dimensional planar slice [splinecam]. SplineCam uses an efficient graph data structure to encode the intersections of the hyperplanes from the various layers that pass through the slice and then uses a fast heuristic breadth-first search algorithm to identify tiles from the graph. All of the computations besides the search can be vectorized and computed on GPUs to enable the visualization of even industrial-scale deep networks.
Figure 5 depicts a SplineCam slice along the plane defined by three training images for a 5-layer ConvNet trained to classify between Egyptian and Tabby cat photos. The first thing we notice is the extraordinarily large number of tiles in just this small region of the 4096-dimensional input space. It can be shown that the decision boundary separating Egyptian and Tabby cats corresponds to a single hyperplane from the final layer that is folded extensively from being pulled back through the previous four layers [romain]. Photos falling in the lower left of the slice are classified as Tabbies, while photos falling in the lower right are classified as Egyptians. The density of tiles also varies across the input space.
@@
It has been known since the late 1980s that even a two-layer neural network is a universal approximator, meaning that, as the number of neurons grows, one can approximate to arbitrary precision an arbitrary continuous function over a Borel measurable set [cybenko]. But, unfortunately, while two-layer networks are easily capable of interpolating a set of training data, in practice they do a poor job generalizing to data outside of the training set. In contrast, deep networks with $L\gg 2$ layers have proved over the past 15 years that they are capable of both interpolating and generalizing well.
@@
Furthermore, deeper networks realize more folded hyperplanes, which can impart more nonlinearity in the spline mapping than a shallower network. For instance, the hyperplane corresponding to a neuron from the first layer can only linearly divide the input space into two half-spaces, while a later-layer neuron can carve up the input space into myriad, even disconnected regions [romain].
Additional work has worked to link the self-similar nature of the tessellation to good approximation. Using self-similarity, one can construct new function spaces for which deeper networks provide better approximation rates (see [devore2021neural, daubechies2022nonlinear] and the references therein). The benefits of depth stem from the fact that the model is able to replicate a part of the function it is trying to approximate in many different places in the input space and with different scalings or orientations. Extending these results, which currently hold only for one-dimensional input and output spaces, to multidimensional signals is an interesting open research avenue. The subdivision results from [romain] could prove useful here.
@@
In particular, we can make a fair, quantitative comparison between the loss landscapes of the ConvNet and ResNet architectures by comparing their singular values. The key finding is that the condition number of a ResNet (the ratio of the largest to smallest singular value) is bounded, while that of the ConvNet is not [rolf]. This means that the local loss landscape of a ResNet with skip connections is provably better conditioned than that of a ConvNet and thus less erratic, less eccentric, and with local minima that are more accommodating to gradient descent optimization.
@@
Astute readers might see a connection to the standard statistical data preprocessing step of data normalization and centering; the main difference is that this processing is performed before each and every gradient learning step. Batch normalization often greatly aids the optimization of a wide variety of deep networks, helping it to find a better (lower) minimum quicker. But the reasons for its efficacy are poorly understood.
We can make progress on understanding batch normalization by again leaning on the affine spline viewpoint. Let’s focus on the effect of batch normalization at initialization just before gradient learning begins; the effect is pronounced, and it is then easy to extrapolate regarding what happens at subsequent gradient steps. Prior to learning, a deep network’s weights are initialized with random values. This means that the initial hyperplane arrangement is also random.
The key finding of [BN-arxiv] is that batch normalization adapts the geometry of a deep network’s spline tessellation to focus the network’s attention on the training data $\left\{\boldsymbol{x}_{i}\right\}_{i=1}^{n}$. It does this by adjusting the angles and offsets of the hyperplanes that form the boundaries of the polytopal tiles to increase their density in regions of the input space inhabited by the training data, thereby enabling finer approximation there.
@@
Figure 8 provides clear evidence of batch normalization’s adaptive prowess. We initialize an 11-layer deep network with a two-dimensional input space three different ways to train on data with a star-shaped distribution. We plot the density of the hyperplanes (basically, the number of hyperplanes passing through local regions of the input space) created by layers 3, 7, and 11 for three different layer configurations: i) the standard layer (6) with bias $\boldsymbol{b}^{(\ell)}={\bf 0}$; ii) the standard layer (6) with random bias $\boldsymbol{b}^{(\ell)}$; iii) the batch normalization layer (8). In all three cases, the weights $\boldsymbol{W}^{(\ell)}$ were initialized randomly. We can make several observations. First, constraining the bias to be zero forces the network into a central hyperplane arrangement tessellation that is incapable of aligning with the data. Second, randomizing both the weights and biases splays the tiles over the entire input space, including many places where the training data is not. Third, batch normalization focuses the hyperplanes from all three of the layers onto the regions where the training data lives.
@@
Current deep learning practice treats a deep network as a black box and optimizes its internal parameters (weights and biases) to minimize some end-to-end training error like the squared loss (4). While this approach has proved mightily successful empirically, it provides no insight into how learning is going on inside the network nor how to improve it. Clearly, as we adjust the parameters to decrease the loss function using gradient descent, the tessellation will change dynamically. Can we use this insight to glean something new about what goes on inside a deep network during learning?
Consider a deep network learning to classify photos of the handwritten digits 0–9. Figure 9 deploys SplineCam to visualize a portion of a 2D slice of the input space of the network defined by three data points in the MNIST handwritten digit training dataset [alwaysgrok]. At left, we see that the tessellation at initialization (before we start learning) is in disarray due to the random weights and biases and nonuse of batch normalization (more on this later). The tessellation is random, and the training error is large.
In the middle, we see the tessellation after convergence to near-zero training error (interpolation), when most of the digits are on the correct side of their respective decision boundaries. Not shown by the figure is the fact that the network also generalizes well to unseen test digits at this juncture. High tile density suggests that even a continuous piecewise affine function can be quite rugged around these points [sebastianParis]. Indeed, the false coloring indicates that the 2-norms of the $\boldsymbol{A}_{\omega}$ matrices has increased around the training images, meaning that their “slopes” have increased. As a consequence, the overall spline mapping $f(\boldsymbol{x})$ is now likely more rugged and more sensitive to changes in the input $\boldsymbol{x}$ as measured by a local (per-tile) Lipschitz constant. In summary, at (near) interpolation, the gradient learning iterations have in some sense accomplished their task (near zero training error) but with elevated sensitivity of $f(\boldsymbol{x})$ to changes in $\boldsymbol{x}$ around the training data points as compared to the random initialization.
@@
Figure 9: SplineCam visualization of a slice of the input space defined by three training MNIST digits being classified by a 4-layer MLP of width 200. The false color map (vivirdis) encodes the 2-norm of the $\boldsymbol{A}_{\omega}$ matrix defined on each tile according to purple (low), green (medium), yellow (high). The decision boundary is depicted in red. (Adapted from [alwaysgrok].)